solve this equation by using the quadratic formula 6x^2=-12x-4

Answer by stanbon(75887)

solve this equation by using the quadratic formula 6x^2=-12x-4
-------------------
6x^2+12x+4 = 0
3x^2+6x+2 = 0
----
x = [-6+- sqrt(36-4*3**2)]/6
----
x = [-6 +- sqrt(12)]6
-----
x = [-6 +-2sqrt(3)]/6
-----
x = [-3 +-sqrt(3)]/3
-----------
Cheers,
Stan H.